An 18 Ω resistor and 30 Ω resistor are connected in series to an unknown voltage source. A very special tool measures the electron flow to be 8.2(10)18 e- to flow through the circuit per second in a clockwise direction.

a. Draw this circuit. Pay close attention to the description!

b. What is the current in the circuit?

c. What is the voltage of the source?

d. What is the total power generated by the components?


1
Expert's answer
2021-04-23T11:11:59-0400

Answer

a. Diagram is as, below




b. Current in circuit

"I=\frac{Q}{t}=\frac{ 8.2\times(10)^{18 }\times1.6\times10^{-19} }{1}\\=1.31A"

c. Voltage

V=IR

"V=1.31\times48=62.88Volts."

d. Total power

P=VI="62.88\times1.31=82.37W"